OpenStudy (anonymous):

MEDAL AND FAN AWARDED Solve the equation. 4 1/5 + b = 9 3/5 a. b=4 4/5 b. b=5 1/5 c. b=5 2/5 d. b=13 4/5 and if you can stay and help with other questions it would be appreciated

OpenStudy (anonymous):

@Luigi0210 you dont know me but i saw you and you are awesome surly you could help me with this?

OpenStudy (anonymous):

thnk you @sammixboo

OpenStudy (anonymous):

can you guys help me?

OpenStudy (luigi0210):

I'm waiting on Sammi :P

OpenStudy (anonymous):

ok

sammixboo (sammixboo):

To solve this you must subtract \(\Large 4\frac{1}{5}\) on both of the equations, like so \(\large 4\frac{1}{5}~ - ~4\frac{1}{5} ~+~b~=~9\frac{3}{5}~-~4\frac{1}{5}\) Since \(\large 4\frac{1}{5}\) minus \(\large 4\frac{1}{5}\) we have 0 + b, which is b So now we have \(\large b~=~9\frac{3}{5}~-~4\frac{1}{5}\) Do you know how to subtract fractions?

OpenStudy (anonymous):

somewhat

sammixboo (sammixboo):

OKy hold on just a second :)

sammixboo (sammixboo):